OpenStudy (anonymous):

What is the Least Common Denominator of 3x/(x+1) + (x+1)/2x + 5/x

OpenStudy (asnaseer):

first list each of the denominators as follows:\[(x+1)\]\[2\times x\]\[x\]then just take all the unique items from this list

OpenStudy (asnaseer):

so you would end up with:\[(x+1)\times2\times x=2x(x+1)\]

OpenStudy (asnaseer):

the last x is ignored because we already had one in 2x

OpenStudy (anonymous):

I Kinda Figured That But I Was 2nd Guessing MySelf :D Thanks

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

I gave you a medal

OpenStudy (asnaseer):

?